The following table contains the discrete probability distribution

      for X= the number of traffic accidents in Norfolk.

X                     0          1          2          3          4          5

P(X)                0.15                 0.10     0.30     0.20     0.08

(a) (4 Points) Determine the missing probability value in the above table.

            Answer: ___________________

(b) (4 points) What is the probability of 2 accidents?

            Answer: ______________________________

(c) (7 points) What is the probability of at least 3 accidents?

            Answer: ________________________

(d) (10 points) Compute the expected value of the number of accidents.
Answer #1

a)

P(1) = 1 - ( 0.15 + 0.1 + 0.3 + 0.2 + 0.08)
= 0.17

b)

P(2) = 0.10

c)
P(at least 3) = 0.3 + 0.2 + 0.08
= 0.58

d)

Expected value = 0 * 0.15 + 1 * 0.17 + 2 * 0.1 + 3 * 0.3 + 4 * 0.2 + 5 * 0.08
= 2.47